Hello. Could anyone please help me with this genetics questions. This is all the information I was given regarding the questions. I would greatly appreciate it.

2. In a herd of 500 cows, each with one calf, weaning weight was recorded as a selection criterion. The mean values for bull and heifer calves were 270Kg and 230 Kg respectively, and the phenotypic standard deviation was P = 40Kg in each sex.

a) What would you expect to be the weight of the heaviest bull (Hint: first determine the proportion selected for one bull, then use the corresponding selection intensity in determining your answer)?

b) If (heritability) of h2 = 0.45, what is the estimated breeding value of this best bull?

c) If he were mated to a randomly selected group of heifers, what would be the expected progeny mean? Assume an equal expression of response in each sex, and an equal sex ratio (50:50).

d) If he were mated to all the heifers of 230Kg and over, what would be the expected progeny mean? (Hint: For a given proportion selected the appendix table does not only contain selection intensity i (average value of selected fraction), but also truncation point x (lowest value of selected fraction).

3. A sheep breeder runs a program to improve wool production in his 1000 ewe flock, and he uses hogget fleece weight as a selection criterion in both sexes. He uses selected rams for two matings and selected ewes for five. Both sexes first drop progeny at 2 years of age, and there is no significant mortality in adults. Each ram is mated to 50 ewes, and the lambing rate is 90%. Assume (heritability) of h2 = 0.55 and (phenotypic Standard Deviation) of P = 0.8Kg. Note: It is handy to summarize the number of males and females used as breeders into a table such as: Age (years) 2 3 4 5 6 Males Females

a) What proportion of males and females must be retained for breeding from each drop or cohort?

b) What is the selection intensity for males and females?

c) What is the predicted response to selection per generation?

d) What is the overall generation interval?

e) What is the predicted response to selection per year?
Answer #1

2.(A) Mean weight of the bulls= 270 kgs with P=40kgs.

on an average the weight of the bulls = 270+/- 40kgs

thus the weight of the heaviest bull= 270+40= 310kgs

(B)heretability =0.45

estimated breeeding value of best bull= 0.45*(270+40), this is the weight of the best bull

answer= 139.5

(C) on breeding of randomly selected heifers the progeny will be that of average weight and equal chances of males and females since there is equal ratio in both sexes.

(D)for heifers of 230kgs and above average weight will be 230+40=270(since P=40 for each sex)

the average weight of progenies will be 230-270kgs